x^2p^²-2xyp+2y^2-x^2=0
1
Expert's answer
2019-12-18T11:39:12-0500

Given equation is :

⟹ x2p2 - 2xyp + 2y2 - x2 = 0


⟹ on solving we get p = ( 2xy + 2x(x2 - y2)1/2 )/2x2


⟹ dy/dx = (y ± (x2 - y2)1/2)/x ..........( 1)

which is homogeneous in x and y

put y=vx so that dy/dx= v+x*dv/dx


∴ from (1), v+dv/dx = (vx± (x2-v2x2)1/2)/x = v ± (1- v)1/2


⟹ x*dv/dx = ±(1 - v2)1/2


⟹ dv/(1 - v2)1/2 = ±dx/x

on integrating we get,

sin-1 v = log x + log c or sin-1 v= -log x - log c

sin-1 y/x = ± log (cx) is the required solution